A drum has certain quantity of pure milk. 12.5 litre of mixture is taken out and replaced by water and the process is repeated two more times. Find initial quantity of milk if in the end the ratio of milk to water is 343:169

Grade:8

1 Answers

Vijay Mukati
8 years ago
Dear Student,

Suppose a container contains x units of liquid from which y units are taken out and replaced by water. then after n operations, the quantity of pure liquid = x(1 - y/x)^n units.

So according to above concept:
Quantity of pure milk left after 3 times replacement will be = x(1-12.5/x)^3 liters
and quanty of water will be = 3 * 12.5 liters

Given ratio of these two is 343:169.
equate the above ratio with this, and get the value of x as your answer.

Thanks.